# Infinite quantum signal processing for arbitrary Szegő functions

### v2

@@ 1.1. Problem setup
then there exists a sequence $\Psi\in\mathbf{P}$ such that $\operatorname{Im}[u_{d}(x,\Psi)]=f(x)$ [31, Theorem 1] (the sequence is not unique in general). QSP only represents polynomials of definite parity. Without loss of generality, we restrict to even functions $f(x)$ throughout the paper, and a similar treatment can be extended to the odd case.
@@
$-\frac{2}{\pi}\int_{0}^{1}\log|1-f(x)^{2}|\frac{dx}{\sqrt{1-x^{2}}}.$
See [30, Lemma 3.1], as well as Section 4.4 below for further details.
@@ 1.2. Main results
(9) $\mathbf{S}_{\eta}=\set{f\in\mathbf{S}}{\left\lVert f\right\rVert_{\infty}\leq 1-\eta}.$
@@
Let $10<\varepsilon<$ 1, $0<\eta<\frac{1}{2}$ and let $k$ and $d$ be integers satisfying $d\geq 1$ and $0\leq k\leq d$. There exists a deterministic algorithm to compute the $k$-th phase factor $\psi_{k}$ for any even input polynomial $f\in\mathbf{S}_{\eta}$ with degree $2d$, to precision $\varepsilon$ with a computational cost of $\mathcal{O}\left(d^{3}+\frac{d\log(d/(\eta\varepsilon))}{\eta}\right)$ and using $\mathcal{O}(\log(d/(\eta\varepsilon)))$ bits. The computational cost of $\mathcal{O}(d^{3})$ arises from solving a linear system of size $\mathcal{O}(d)$. To determine all phase factors, the algorithm solves $\mathcal{O}(d)$ such linear systems, resulting in a cumulative cost of $\mathcal{O}\left(d^{4}+\frac{d\log(d/(\eta\varepsilon))}{\eta}\right)$, and the bit requirement remains $\mathcal{O}(\log(d/(\eta\varepsilon)))$.
@@ 1.4. Discussion and open questions
Our algorithm in Theorem 2 allows us to compute all phase factors independently and in parallel. This completely circumvents the error accumulation issue associated with the “layer stripping” method introduced in [13], which can be another source of numerical instability and requires at least in theory, high precision arithmetics [15, Eq. (38)]. The computation of each phase factor requires solving a Riemann-Hilbert factorization problem via a linear system of equations, leading to a $\mathcal{O}(d^{3})$ cost per phase factor. Consequently, the leading term of the computational cost in Theorem 2 is $\mathcal{O}(d^{4})$. On the other hand, the FPI algorithm [8] only requires a much lower cost of $\mathcal{O}(d^{2}\log(1/\varepsilon))$ upon convergence. The lower bound for evaluating phase factors for all functions $f\in\mathbf{S}$ is unclear either at this point.
Ref. [10] observed a close connection between the decay of the Chebyshev coefficients of $f$ and the decay of the phase factors. This decay relationship was rigorously established under $\ell^{1}$ conditions in [8, Theorem 4]. We anticipate that this relationship will generalize to all functions $f\in\mathbf{S}$, which will be our future work.
@@ Acknowledgment
Note: Towards the completion of this work, we learned that Hongkang Ni is developing a new numerically stable method based on our findings, which may compute all phase factors with a cost of $\widetilde{\mathcal{O}}(d^{2}\log(1/\varepsilon))$.
@@ 2.1. Hardy functions
Functions in $H^{p}(\mathbb{D})$ have radial limits almost everywhere on the unit circle $\mathbb{T}:=\{z\in\mathbb{C}:\left\lvert z\right\rvert=1\}$, and these boundary values determine the function uniquely. Thus when we say a function $g\in L^{p}(\mathbb{T})$ belongs to $H^{p}(\mathbb{D})$, we mean $g$ coincides with the boundary values of a unique $H^{p}(\mathbb{D})$ function, which we also denote by $g$. By the mean value property for harmonic functions, for every function $g\in H^{p}(\mathbb{D})$ we have
[…]
@@
More generally, [12, Chapter 3, Theorem 2.1 and Theorem 2.3] shows that $\mathrm{H}$ extends to a bounded operator on $L^{p}(\mathbb{T})$ for $1<p<\infty$, and is a bounded map from $L^{1}(\mathbb{T})$ to $L^{1,\infty}(\mathbb{T})$. Given $g\in H^{p}(\mathbb{D})$, let us denote $g(e^{i\theta})=\sum_{n\geq 0}c_{n}e^{in\theta}$. If $c_{0}=0$, then
@@
If $g\in H^{p}(\mathbb{D})$ has modulus $1$ a.e. on $\mathbb{T}$, then $g$ is called an inner function. Two functions $a,b\in H^{2}(\mathbb{D})$ are said to have no common inner factor if for every inner function $g$, both $a/g$ and $b/g$ are $H^{2}(\mathbb{D})$ functions if and only if $g$ is constant. A function $g\in L^{\infty}(\mathbb{T})$ is called outer if $\log|g|\in L^{1}(\mathbb{T})$ and $g=e^{G}$ where $G=\log|g|+i\mathrm{H}(\log|g|)$. Note that if $g$ is outer, then $G$ has an analytic extension to $\mathbb{D}$ with real part bounded above, and hence $g\in H^{\infty}(\mathbb{D})$.
@@ 2.2. Nonlinear Fourier analysis
$\mathbf{H}_{\leq k}:=\{(a,b)\leavevmode\nobreak\ :\leavevmode\nobreak\ (a,b^{*})\in\mathbf{H}_{\geq-k}\}\,,$
@@
One can also extend the nonlinear Fourier transform to sequences in $\ell^{2}(\mathbb{Z})$ as follows: given a sequence $F$ in $\ell^{2}(\mathbb{Z})$, split it as $F_{-}+F_{+}$, where $F_{-}$ is supported in $(-\infty,-1]$ and $F_{+}$ is supported in $[0,\infty)$. Then we define the nonlinear Fourier transform of $F$ to be the pair
@@
where $(a_{-},b_{-})$ denotes the NLFT of $F_{-}$ and $(a_{+},b_{+})$ denotes the NLFT of $F_{+}$. The problem of finding factors $(a_{-},b_{-})$ and $(a_{+},b_{+})$ as in Eq. 25 is known as a Riemann-Hilbert factorization problem [27, Lecture 3, p.31].
@@ 2.3. Weiss algorithm for constructing $a$ given $b$
It follows that
$a^{*}(0)=e^{G^{*}(0)}>0\,,$
[…]
@@ 3.1. Outline
since this is a sufficient and necessary condition for the existence of QSP representation for polynomial $f$ with definite parity [13, Corollary 5]. We provide a positive answer in Theorem 5 below, extending [1, Theorem 11].
@@ 3.2. Weiss algorithm
The construction of $a$ as shown in the proof of Theorem 4 consists of three steps: $R(z):=\log\sqrt{1-\left\lvert b(z)\right\rvert^{2}}$; $G(z):=R(z)-\mathrm{i}\mathrm{H}(R(z))$; and $a(z):=e^{G(z)}$. To construct $\frac{b}{a}$, we only need to replace the last step by evaluating $\frac{b}{a}:=b(z)e^{-G(z)}$. In our method, we directly evaluate the Fourier coefficients of $\frac{b}{a}$ with indices ranging from 00 to $d$ using the Fast Fourier Transform (FFT). Let $b(z)$ be a Laurent polynomial of degree $d$. Due to the correspondence between the target function $f(x)$ and $b(z)$, we have $b(z)=b(z^{-1})$ and $\left\lVert b\right\rVert_{\infty}=\left\lVert f\right\rVert_{\infty}\leq 1-\eta$. We also denote by $N$ the discretization parameter associated with FFT.
@@ 3.3. Riemann-Hilbert factorization
belongs to $H^{2}(\mathbb{D})$ and has real Fourier coefficients as well. Since the Fourier expansion of $G$ coincides with its Taylor expansion as a holomorphic function in the unit disk $\mathbb{D}$, then $G(z)$ and all of its derivatives, when evaluated at $z=0$, must be real-valued. Thus the $H^{\infty}(\mathbb{D})$ function
@@
Given any $0<\eta<\frac{1}{2}$, $10<\varepsilon<$ 1 and integer $d\geq 1$, assume $f\in\mathbf{S}_{\eta}$ to be a polynomial of degree $2d$. Let
@@ 3.4. Numerical experiment
(a) The magnitude of phase factors $\Psi$ for each index $i$.
@@
(a) The magnitude of phase factors $\Psi$ for each index $i$.
@@
In this section, we present the proof of Theorem 1 which provides a representation of measurable functions by a square summable sequence. As mentioned previously in Section 2.4 and Section 3.1, we prove Theorem 5, extending the Riemann-Hilbert factorization result [1, Theorem 11] in nonlinear Fourier analysis to all $(a,b)\in\mathbf{B}$. Our key idea here is to argue that the operator $M$ arising in Eq. 39 is antisymmetric on the Hilbert space $H^{2}(\mathbb{D}^{*})\times H^{2}(\mathbb{D})$, meaning it possesses pure imaginary spectrum and hence $\operatorname{Id}+M$ is invertible. If $(a,b)\in\mathbf{B}$ satisfies $|a|>\eta^{\prime}$ for some $\eta^{\prime}>0$, then $\frac{b}{a}$ is bounded on $\mathbb{T}$ and the above reasoning is rigorous. However, if $(a,b)\in\mathbf{B}$, then $\frac{b}{a}$ may not be integrable and $M$ may not be well-defined on $H^{2}(\mathbb{D}^{*})\times H^{2}(\mathbb{D})$. This leads to the technicalities in this section as we must appropriately interpret $M$ at the endpoint case $\eta^{\prime}=0$. Indeed, the simple operator $M$ in Eq. 39 must be interpreted as a densely defined operator as in Eq. 62, which involves truncating $a$ from below and taking a weak limit the in the truncation parameter. And it’s by using the theory of unbounded operators [26, Chapter 13] that we argue that $M$ is an antisymmetric unbounded operator to then get $\operatorname{Id}+M$ has bounded inverse. The arguments in the case $|a|>\eta^{\prime}$ for some $\eta^{\prime}>0$, which follows from $f\in\mathbf{S}_{\eta}$ for some $\eta>0$, are considerably simpler as the weak limits coincide with the simple operator in Eq. 39. For instance, Lemma 10 is near immediate if we take $M$ as in Eq. 39. Thus any reader uninterested in the endpoint case $\eta=0$ may read this section while ignoring all weak limits.
@@ 4.1. Inversion of an unbounded operator using spectral theory
[…]
@@
The space $\mathcal{H}$ is closed. The space $\mathcal{D}$ is dense in $\mathcal{H}$ by Beurling’s theorem [12, Chapter 2, Corollary 7.3], which says that the set of all polynomials in $z$, when multiplied by an outer function on $\mathbb{D}$, is dense in $H^{2}(\mathbb{D})$.
@@
If $|a|$ is bounded below uniformly away from 00, then $\mathcal{E}=\mathcal{H}$ and for all $\eta$ sufficiently small Eq. 58 equals
@@
$=\left\|\begin{pmatrix}0&b^{*}\frac{a^{*}}{a^{*}_{\eta}}\\ -b\frac{a}{a_{\eta}}&0\end{pmatrix}h-\begin{pmatrix}0&b^{*}\\ -b&0\end{pmatrix}h\right\|=\left\|\begin{pmatrix}0&b^{*}\\ -b&0\end{pmatrix}\begin{pmatrix}\frac{a}{a_{\eta}}-1&0\\ 0&\frac{a^{*}}{a^{*}_{\eta}}-1\end{pmatrix}h\right\|\,.$
@@
We refer to [26, Chapter 13] for the spectral theory of unbounded operators on a Hilbert space.
@@
Recall from [26, Definition 13.1] that given the densely defined operator $M:\mathcal{D}(M)\to\mathcal{H}$, the domain $\mathcal{D}(M^{*})$ of the adjoint operator $M^{*}$ consists of all those $g\in\mathcal{H}$ for which
@@
To show $\mathrm{i}M$ is self-adjoint [26, Definition 13.3], we must check that for all $f,g\in\mathcal{D}(M)$, we have
@@
(68) $-\langle\begin{pmatrix}0&\frac{b^{*}}{a_{\eta}^{*}}\\ -\frac{b}{a_{\eta}}&0\end{pmatrix}g,h\rangle=\langle g,\begin{pmatrix}0&\frac{b^{*}}{a_{\eta}^{*}}\\ -\frac{b}{a_{\eta}}&0\end{pmatrix}h\rangle=\langle g,\mathcal{P}_{\mathcal{H}}\begin{pmatrix}0&\frac{b^{*}}{a_{\eta}^{*}}\\ -\frac{b}{a_{\eta}}&0\end{pmatrix}h\rangle\,,$
@@
Since $\mathrm{i}M$ is self-adjoint, then by [26, Theorem 13.30] it has real spectrum. ∎
By Lemma 10, $M$ has pure imaginary spectrum, and so for any nonzero real $\lambda$, the operator $\operatorname{Id}+\lambda M$ is a linear bijection from $\mathcal{E}$ onto $\mathcal{H}$ with bounded inverse [26, Definition 13.26]. We also have the operator norm estimate
@@
$+\langle\lambda Mx,x\rangle+\langle x,\lambda Mx\rangle=\|x\|^{2}+\|\lambda Mx\|^{2}\geq\|x\|^{2}\,.$

### v3

@@ 1.1. Problem setup
then there exists a sequence $\Psi\in\mathbf{P}$ such that $\operatorname{Im}[u_{d}(x,\Psi)]=f(x)$ [29, Theorem 1] (the sequence is not unique in general). QSP only represents polynomials of definite parity. Without loss of generality, we restrict to even functions $f(x)$ throughout the paper, and a similar treatment can be extended to the odd case.
@@
$\frac{2}{\pi}\int_{0}^{1}\log|1-f(x)^{2}|\frac{dx}{\sqrt{1-x^{2}}}.$
See [28, Lemma 3.1], as well as Section 4.4 below for further details.
@@ 1.2. Main results
[…]
@@
Let $0<\varepsilon<1$, $0<\eta<\frac{1}{2}$ and let $k$ and $d$ be integers satisfying $d\geq 1$ and $0\leq k\leq d$. There exists a deterministic algorithm to compute the $k$-th phase factor $\psi_{k}$ for any even input polynomial $f\in\mathbf{S}_{\eta}$ with degree $2d$, to precision $\varepsilon$ with a computational cost of $\mathcal{O}\left(d^{3}+\frac{d}{\eta}\log^{2}(d/(\eta\varepsilon))\right)$ and using $\mathcal{O}(\log(d/(\eta\varepsilon)))$ bits. The computational cost of $\mathcal{O}(d^{3})$ arises from solving a linear system of size $\mathcal{O}(d)$. To determine all phase factors, the algorithm solves $\mathcal{O}(d)$ such linear systems, resulting in a cumulative cost of $\mathcal{O}\left(d^{4}+\frac{d}{\eta}\log^{2}(d/(\eta\varepsilon))\right)$, and the bit requirement remains $\mathcal{O}(\log(d/(\eta\varepsilon)))$.
@@ 1.4. Discussion and open questions
Our algorithm in Theorem 2 allows us to compute all phase factors independently and in parallel. This completely circumvents the error accumulation issue associated with the “layer stripping” method introduced in [12], which can be another source of numerical instability and requires at least in theory, high precision arithmetics [14, Eq. (38)]. The computation of each phase factor requires solving a Riemann-Hilbert factorization problem via a linear system of equations, leading to a $\mathcal{O}(d^{3})$ cost per phase factor. Consequently, the leading term of the computational cost in Theorem 2 is $\mathcal{O}(d^{4})$. On the other hand, the FPI algorithm [8] only requires a much lower cost of $\mathcal{O}(d^{2}\log(1/\varepsilon))$ upon convergence. The lower bound for evaluating phase factors for all functions $f\in\mathbf{S}$ is unclear either at this point.
Ref. [10] observed a close connection between the decay of the Chebyshev coefficients of $f$ and the decay of the phase factors. This decay relationship was rigorously established under $\ell^{1}$ conditions in [8, Theorem 4]. We anticipate that this relationship may be generalized to all functions $f\in\mathbf{S}$.
@@ Acknowledgment
Note: Towards the completion of this work, we learned that Hongkang Ni and Lexing Ying were developing a new numerically stable method based on our findings, which can reduce the cost of computing all phase factors to $\widetilde{\mathcal{O}}(d^{2}\log(1/\varepsilon))$.
@@ 2.1. Hardy functions
Functions in $H^{p}(\mathbb{D})$ have radial limits almost everywhere on the unit circle $\mathbb{T}:=\{z\in\mathbb{C}:\left\lvert z\right\rvert=1\}$, and these boundary values determine the function uniquely [11, Chapter 2, Thm. 3.1 and Cor. 3.2]. Thus when we say a function $g\in L^{p}(\mathbb{T})$ belongs to $H^{p}(\mathbb{D})$, we mean $g$ coincides with the boundary values of a unique $H^{p}(\mathbb{D})$ function, which we also denote by $g$. By the mean value property for harmonic functions, for every function $g\in H^{p}(\mathbb{D})$ we have
$g(0)=\int\limits_{\mathbb{T}}g:=\frac{1}{2\pi}\int_{0}^{2\pi}g(e^{i\theta})d\theta\,.$
@@
More generally, [11, Chapter 3, Theorem 2.1 and Theorem 2.3] shows that $\mathrm{H}$ extends to a bounded operator on $L^{p}(\mathbb{T})$ for $1<p<\infty$, and is a bounded map from $L^{1}(\mathbb{T})$ to $L^{1,\infty}(\mathbb{T})$. Given $g\in H^{p}(\mathbb{D})$, let us denote $g(e^{i\theta})=\sum_{n\geq 0}c_{n}e^{in\theta}$. If $c_{0}=0$, then
@@
If $g\in H^{p}(\mathbb{D})$ has modulus $1$ a.e. on $\mathbb{T}$, then $g$ is called an inner function. Two functions $a,b\in H^{2}(\mathbb{D})$ are said to have no common inner factor if for every inner function $g$, both $a/g$ and $b/g$ are $H^{2}(\mathbb{D})$ functions if and only if $g$ is constant. A function $g\in L^{\infty}(\mathbb{T})$ is called outer if $\log|g|\in L^{1}(\mathbb{T})$ and $g=e^{G}$ where $G=\log|g|+i\mathrm{H}(\log|g|)$. Note that if $g$ is outer, then $G$ has an analytic extension to $\mathbb{D}$ [11, first page of Chapter 3], whose real part is at most $\sup\limits_{z\in\mathbb{T}}\log|g(z)|$, and hence $g\in H^{\infty}(\mathbb{D})$.
@@ 2.2. Nonlinear Fourier analysis
$\mathbf{H}_{\leq k}:=\{(a,b)\penalty 10000\ :\penalty 10000\ (a,b^{*})\in\mathbf{H}_{\geq-k}\}\,$,
@@
One can also extend the nonlinear Fourier transform to sequences in $\ell^{2}(\mathbb{Z})$ as follows: given a sequence $F$ in $\ell^{2}(\mathbb{Z})$, split it as $F_{-}+F_{+}$, where $F_{-}$ is supported in $(-\infty,-1]$ and $F_{+}$ is supported in $0,\infty)$. Then we define the nonlinear Fourier transform of $F$ to be the pair
@@
where $(a_{-},b_{-})$ denotes the NLFT of $F_{-}$ and $(a_{+},b_{+})$ denotes the NLFT of $F_{+}$. The problem of finding factors $(a_{-},b_{-})$ and $(a_{+},b_{+})$ as in Eq. 25 is known as a Riemann-Hilbert factorization problem [26, Lecture 3, p.31].
@@ 2.3. Weiss algorithm for constructing $a$ given $b$
Note that since conjugate functions vanish at the origin, then $G^{*}$ has vanishing imaginary part at the origin [11, first page of Chapter 3]. It follows that $G^{*}(0)$ is the mean of $R=\log\sqrt{1-|b(z)|^{2}}$ on $\mathbb{T}$, which is real-valued and finite by $\|b\|_{\infty}\leq 1$ and the Szegő condition Eq. 34. It follows that
$a^{*}(0)=e^{G^{*}(0)}>0\,.$
@@ 3.1. Outline
since this is a sufficient and necessary condition for the existence of QSP representation for polynomial $f$ with definite parity [12, Corollary 5]. We provide a positive answer in Theorem 5 below, extending [1, Theorem 11].
@@ 3.2. Weiss algorithm
The construction of $a$ as shown in the proof of Theorem 4 consists of three steps: $R(z):=\log\sqrt{1-\left\lvert b(z)\right\rvert^{2}}$; $G(z):=R(z)-\mathrm{i}\mathrm{H}(R(z))$; and $a(z):=e^{G(z)}$. To construct $\frac{b}{a}$, we only need to replace the last step by evaluating $\frac{b}{a}:=b(z)e^{-G(z)}$. In our method, we directly evaluate the Fourier coefficients of $\frac{b}{a}$ with indices ranging from 0 to $d$ using the Fast Fourier Transform (FFT). Let $b(z)$ be a Laurent polynomial of degree $d$. Due to the correspondence between the target function $f(x)$ and $b(z)$, we have $b(z)=b(z^{-1})$ and $\left\lVert b\right\rVert_{\infty}=\left\lVert f\right\rVert_{\infty}\leq 1-\eta$. We also denote by $N$ the discretization parameter associated with FFT.
@@ 3.3. Riemann-Hilbert factorization
belongs to $H^{2}(\mathbb{D})$ and has real Fourier coefficients as well, since by (16) the operator $\mathrm{i}\mathrm{H}$ is a real-valued multiplier on the frequency side. Since the Fourier expansion of $G$ coincides with its Taylor expansion as a holomorphic function in the unit disk $\mathbb{D}$, then $G(z)$ and all of its derivatives, when evaluated at $z=0$, must be real-valued. Thus the $H^{\infty}(\mathbb{D})$ function
@@
Given any $0<\eta<\frac{1}{2}$, $0<\varepsilon<1$ and integer $d\geq 1$, assume $f\in\mathbf{S}_{\eta}$ to be a polynomial of degree $2d$. Let
@@ 3.4. Numerical experiment
(a) The magnitude of phase factors $\Psi$ for each index ii.
@@
(a) The magnitude of phase factors $\Psi$ for each index ii.
@@
In this section, we present the proof of Theorem 1 which provides a representation of measurable functions by a square summable sequence. As mentioned previously in Section 2.4 and Section 3.1, we prove Theorem 5, extending the Riemann-Hilbert factorization result [1, Theorem 11] in nonlinear Fourier analysis to all $(a,b)\in\mathbf{B}$. Our key idea here is to argue that the operator $M$ arising in Eq. 39 is antisymmetric on the Hilbert space $H^{2}(\mathbb{D}^{*})\times H^{2}(\mathbb{D})$, meaning it possesses pure imaginary spectrum and hence $\operatorname{Id}+M$ is invertible. If $(a,b)\in\mathbf{B}$ satisfies $|a|>\eta^{\prime}$ for some $\eta^{\prime}>0$, then $\frac{b}{a}$ is bounded on $\mathbb{T}$ and the above reasoning is rigorous. However, if $(a,b)\in\mathbf{B}$, then $\frac{b}{a}$ may not be integrable and $M$ may not be well-defined on $H^{2}(\mathbb{D}^{*})\times H^{2}(\mathbb{D})$. This leads to the technicalities in this section as we must appropriately interpret $M$ at the endpoint case $\eta^{\prime}=0$. Indeed, the simple operator $M$ in Eq. 39 must be interpreted as a densely defined operator as in Eq. 62, which involves truncating $a$ from below and taking a weak limit the in the truncation parameter. And it’s by using the theory of unbounded operators [25, Chapter 13] that we argue that $M$ is an antisymmetric unbounded operator to then get $\operatorname{Id}+M$ has bounded inverse. The arguments in the case $|a|>\eta^{\prime}$ for some $\eta^{\prime}>0$, which follows from $f\in\mathbf{S}_{\eta}$ for some $\eta>0$, are considerably simpler as the weak limits coincide with the simple operator in Eq. 39. For instance, Lemma 10 is near immediate if we take $M$ as in Eq. 39. Thus any reader uninterested in the endpoint case $\eta=0$ may read this section while ignoring all weak limits.
@@ 4.1. Inversion of an unbounded operator using spectral theory
induced by the standard inner product $\langle\cdot,\cdot\rangle$ on $L^{2}(\mathbb{T})\times L^{2}(\mathbb{T})$. As $a^{*}$ is bounded and outer, we have
@@
The space $\mathcal{H}$ is closed. The space $\mathcal{D}$ is dense in $\mathcal{H}$ by Beurling’s theorem [11, Chapter 2, Corollary 7.3], which says that the set of all polynomials in $z$, when multiplied by an outer function on $\mathbb{D}$, is dense in $H^{2}(\mathbb{D})$.
@@
If $|a|$ is bounded below uniformly away from 0, then $\mathcal{E}=\mathcal{H}$ and for all $\eta$ sufficiently small Eq. 58 equals
@@
$\left\|\begin{pmatrix}0&b^{*}\frac{a^{*}}{a^{*}_{\eta}}\\ -b\frac{a}{a_{\eta}}&0\end{pmatrix}h-\begin{pmatrix}0&b^{*}\\ -b&0\end{pmatrix}h\right\|=\left\|\begin{pmatrix}0&b^{*}\\ -b&0\end{pmatrix}\begin{pmatrix}\frac{a}{a_{\eta}}-1&0\\ 0&\frac{a^{*}}{a^{*}_{\eta}}-1\end{pmatrix}h\right\|\,.$
@@
We refer to [25, Chapter 13] for the spectral theory of unbounded operators on a Hilbert space.
@@
Recall from [25, Definition 13.1] that given the densely defined operator $M:\mathcal{D}(M)\to\mathcal{H}$, the domain $\mathcal{D}(M^{*})$ of the adjoint operator $M^{*}$ consists of all those $g\in\mathcal{H}$ for which
@@
To show $\mathrm{i}M$ is self-adjoint [25, Definition 13.3], we must check that for all $f,g\in\mathcal{D}(M)$, we have
@@
(68) $\langle\begin{pmatrix}0&\frac{b^{*}}{a_{\eta}^{*}}\\ -\frac{b}{a_{\eta}}&0\end{pmatrix}g,h\rangle=\langle g,\begin{pmatrix}0&\frac{b^{*}}{a_{\eta}^{*}}\\ -\frac{b}{a_{\eta}}&0\end{pmatrix}h\rangle=\langle g,\mathcal{P}_{\mathcal{H}}\begin{pmatrix}0&\frac{b^{*}}{a_{\eta}^{*}}\\ -\frac{b}{a_{\eta}}&0\end{pmatrix}h\rangle\,$,
@@
Since $\mathrm{i}M$ is self-adjoint, then by [25, Theorem 13.30] it has real spectrum. ∎
By Lemma 10, $M$ has pure imaginary spectrum, and so for any nonzero real $\lambda$, the operator $\operatorname{Id}+\lambda M$ is a linear bijection from $\mathcal{E}$ onto $\mathcal{H}$ with bounded inverse [25, Definition 13.26]. We also have the operator norm estimate
@@
$\langle\lambda Mx,x\rangle+\langle x,\lambda Mx\rangle=\|x\|^{2}+\|\lambda Mx\|^{2}\geq\|x\|^{2}\,.$